# Stochastic Bregman Proximal Gradient Method Revisited: Kernel Conditioning and Painless Variance Reduction

### v3

@@ 2.5 A Lipschitz-like bound for gradient difference
Even with a single sample $\xi$, the above MSE will automatically go to 00 if $\|x-x_{\text{ref}}\|\to 0$. For the variance reduced methods whose design ensures this to happen, given an accurate enough estimator ~$\tilde{\nabla}f(x_{\text{ref}})\approx\nabla f(x_{\text{ref}})$, then ~$\tilde{\nabla}f(x_{\text{ref}})+\Delta_{\xi}$ will give a much more accurate estimation of $\nabla f(x)$. This forms the basic insight why variance reduction accelerates SGD under classic $L$-smoothness condition.
@@
[…]
@@ 3.1 The general algorithm and analysis framework
1 Input: Initial point $x_{1,0}$, constant $\delta$, epoch length $\tau$, step size $\eta$, interpolation factor $\gamma\in(0,1]$.
@@
3       Construct a convex set $\mathcal{X}_{s}\supseteq{B}(x_{s,0},\delta/2)$ such that $\kappa_{h}(\mathcal{X}_{s})\leq\kappa_{h}^{\delta}$
@@ 7             If $k\geq 1$, uniformly sample a mini-batch $\mathcal{B}_{s,k}\subseteq[n]$ with replacement, compute
$\bar{x}_{s,k+1}=\mathbf{T}_{\phi+\mathrm{id}_{\mathcal{X}_{s}}\!,h}^{\eta}(x_{s,k},v_{s,k})\qquad\mbox{and}\qquad x_{s,k+1}=(1-\gamma)x_{s,k}+\gamma\bar{x}_{s,k+1}\vspace{-0.1cm}$ (18)
@@
Using the fact that $L^{2}=\frac{1}{n}\sum_{i=1}^{n}L_{i}^{2}$ and the fact that $\kappa_{h}(\mathcal{X}_{s})\leq\kappa_{h}^{\delta}$ leads to (22).
@@ 3.2.2 General non-element-separable cases
For the linear consensus constraint $x-y=0$, the corresponding coefficient matrix is $[I_{d\times d},-I_{d\times d}]$ and its condition number is 1. Then many primal-dual algorithms can achieve an $O(\kappa_{h}^{\delta}\ln\epsilon_{\mathrm{tol}}^{-1})$ complexity for finding an $\epsilon_{\mathrm{tol}}$-optimal solution, see [49]. Due to Lemma 3.9, we only need to deal with this scenario for limited times. This is also how we implement the subproblem solvers in the experiments.
@@ 3.2.3 Composition of norm cases
Computing the objective gradient of (40) gives $H^{\prime}(\alpha)-\|u\|$. Because $H(\cdot)$ is strictly convex and increasing, we know $H^{\prime}(\cdot)\geq 0$ and $H^{\prime}(\cdot)$ is increasing on $[0,+\infty)$, then we know $\alpha^{*}=0$ if $H^{\prime}(0)\geq\|u\|$. Otherwise, there is a unique solution $\alpha^{*}$ s.t. $H^{\prime}(\alpha^{*})=\|u\|$. In some cases, the equation $H^{\prime}(\alpha)=\|u\|$ allows a closed form solution, then we can directly adopt it, see [5]. If no closed-form solution is available, then one can apply either Newton’s method or a binary search to find the root. Based on this observation, let us consider a few more examples.
@@
In fact, this result can be generalized to the group $\ell_{1}$/$\ell_{2}$ norm that promotes group sparsity.
@@ 4.1 An adaptive step size control for BPG method
[…]
@@ 4.2 Adaptive step size control with stochastic variance reduction
$\bar{x}_{s,k+1}=\mathbf{T}_{\phi,\,h}^{\eta_{s,k}}\big{(}x_{s,k},v_{s,k}\big{)}\qquad\mbox{with}\qquad\eta_{s,k}=\min\left\{\frac{1}{2\kappa_{h}^{\delta}L},\,\frac{\mu\delta}{3\rho},\,\frac{\mu\delta}{\|v_{s,k}\|+\rho}\right\}\,,\quad$ (46)
@@
Suppose $(E,\left\lvert\left\lvert\left\lvert\cdot\right\rvert\right\rvert\right\rvert)$ is $\kappa$-regular for some $\kappa\geq 1$ and $\{\zeta_{t}\}_{t\geq 0}$ is an $E$-valued martingale difference sequence w.r.t. the filtration $\{\mathcal{F}_{t}\}_{t\geq 0}$ and default $\zeta_{0}=0$. Suppose $\{\zeta_{t}\}_{t\geq 0}$ satisfies the following light-tail property:
@@
Suppose Assumptions 2.9 and 3.1 hold. For any epoch $s\geq 1$ and let $p_{s}=\frac{6q}{\pi^{2}s^{2}\tau}$ for some $q\in(0,1)$. Suppose we select $|\mathcal{B}_{s,k}|=b_{s}$ for all $10\leq k\leq\tau-$ 1, then
@@
In the above definition, the index $k$ runs through $11\leq k\leq\tau-$ 1, and the index $j$ can take value from $1\leq j\leq|\mathcal{B}_{s,k}|$ given each $k$. For each $(k,j)$ in our index range, $\xi_{k,j}^{s}$ stands for the $j$-th sample from the batch $\mathcal{B}_{s,k}\subseteq[n]$. Then by direct computation, we have $\mathcal{E}_{s,k}=\sum_{k^{\prime}=1}^{k}\sum_{j=1}^{|\mathcal{B}_{s,k^{\prime}}|}\zeta_{k^{\prime},j}^{s}$ and $\{\zeta_{k,j}^{s}\}$ forms a martingale difference sequence if the index $(k,j)$ runs in a lexicographical order. Note that
@@
Let us set the parameters according to Lemma 4.9 and 4.10. Consider the target accuracy $\epsilon\leq\max\left\{2L\kappa_{h}^{\delta}\mu\delta\rho/3,\rho^{2}\right\}$, let $\mathcal{X}_{\epsilon}:=\mathrm{Lev}_{0}\cap B(x_{1,0},R_{\epsilon})$ be a compact set, then with probability at least $1-q$, all the following arguments hold:
@@
Moreover, we have $\|\mathcal{D}_{\phi,h}^{\eta_{S_{\epsilon},K_{\epsilon}}}(x_{S_{\epsilon},K_{\epsilon}})\|^{2}\leq 2.5\epsilon$. The total samples consumed is $\sum_{s=1}^{S_{\epsilon}}n+\tau b_{s}=\widetilde{O}(\sqrt{n}/\epsilon)$, where $\widetilde{O}(\cdot)$ hides the poly-logarithmic terms in $q$ and $\epsilon$.
@@ 5 Numerical experiments
Figure 4: Experimental results for quadratic reverse problem. The raw signal vector $x_{\mathrm{true}}$ is attached at the bottom left corner of each subfigure. The “# of datapass” in x-axis refers to #samples consumedfull batch size $\frac{\#\mbox{samples consumed}}{\mbox{full batch size }N}$.
@@ Appendix B Proof of Section 2
First, direct computation gives $\nabla^{2}h(x)=(\|x\|^{r}+{\color[rgb]{0,0,0}\alpha})\cdot I+r\|x\|^{r-2}\cdot xx^{\top}$. For $\forall x\in\mathbb{R}^{d}$, we have $\lambda_{\max}(\nabla^{2}h(x))=(r+1)\|x\|^{r}+{\color[rgb]{0,0,0}\alpha}$ and $\lambda_{\min}(\nabla^{2}h(x))=\|x\|^{r}+{\color[rgb]{0,0,0}\alpha}$. Then for any compact set $\mathcal{X}\subseteq\mathbb{R}^{n}$ with diameter denoted by $\mathrm{diam}(\mathcal{X})=d_{\mathcal{X}}$, let $y\in\mathop{\mathrm{argmax}}_{u\in\mathcal{X}}\|u\|$ and $x\in\mathop{\mathrm{argmin}}_{u\in\mathcal{X}}\|u\|$, then
$\kappa_{h}(\mathcal{X})=\frac{\lambda_{\max}(\nabla^{2}h(y))}{\lambda_{\min}(\nabla^{2}h(x))}=\frac{(r+1)\|y\|^{r}+{\color[rgb]{0,0,0}\alpha}}{\|x\|^{r}+{\color[rgb]{0,0,0}\alpha}}\leq\frac{(r+1)(\|x\|+{\color[rgb]{0,0,0}d_{\mathcal{X}}})^{r}+{\color[rgb]{0,0,0}\alpha}}{\|x\|^{r}+{\color[rgb]{0,0,0}\alpha}}.$ (52)
@@
$\kappa_{h}(\mathcal{X})\leq(r+1)\left(1+\frac{d_{\mathcal{X}}}{\|x\|}\right)^{r}+\frac{{\color[rgb]{0,0,0}\alpha}}{\|x\|^{r}+{\color[rgb]{0,0,0}\alpha}}\leq(r+1)\left(1+\frac{1}{r}\right)^{r}+1\leq 3r+4,\quad\forall r\geq 0.$ (53)
@@
$\kappa_{h}^{\delta}\leq(r+1)\cdot\sup\left\{\frac{(t+\delta)^{r}}{t^{r}+{\color[rgb]{0,0,0}\alpha}}:t\geq 0\right\}+1.$ (54)
By direct computation, the function $\rho(t):=\frac{(t+\delta)^{r}}{t^{r}+{\color[rgb]{0,0,0}\alpha}}$ attains its maximal value at
$t^{*}=\begin{cases}\alpha^{\frac{1}{r^{2}-r}}\delta^{-\frac{1}{r-1}},&\mbox{ if }r>1\\ 0,&\mbox{ if }r\leq 1,\delta^{r}\geq\alpha\\ +\infty,&\mbox{ if }r\leq 1,\delta^{r}<\alpha\end{cases}\qquad\mbox{with}\qquad\rho(t^{*})=\begin{cases}\big{(}1+\delta^{\frac{r}{r-1}}\alpha^{-\frac{1}{r-1}}\big{)}^{r-1},&\mbox{ if }r>1\\ \delta^{r}/\alpha,&\mbox{ if }r\leq 1,\delta^{r}\geq\alpha\\ 1,&\mbox{ if }r\leq 1,\delta^{r}<\alpha\end{cases}$
@@
In particular, when $r>1$ and $\delta\leq\alpha^{\frac{1}{r}}/r\leq\alpha^{\frac{1}{r}}/r^{\frac{r-1}{r}}$, we have $\big{(}\frac{\delta^{r}}{\alpha}\big{)}^{\frac{1}{r-1}}\leq\frac{1}{r-1}$. Following the same logic of the last inequality in (53), we obtain $\kappa_{h}^{\delta}\leq 3r+4$ in this case. This completes the proof.
@@
Let $A$ be a symmetric matrix, and let $B\succeq 0$ be positive semidefinite matrix. Then $-B\preceq A\preceq B$ indicates that $\|A\|\leq\|B\|$.
@@
As $\mathcal{X}$ is convex, the line segment $[x,y]\subseteq\mathcal{X}$ and hence
@@ Appendix E Extension to multi-block kernel conditioning
Finally, we will roughly discuss how one can extend KC-regularity to block separable kernels, which is considered in [3, 15, 16, 19]. In particular, we will mostly focus on the extension of Section 3, while the extension of Section 4 can be done similarly. To differentiate the meaning of the subscripts, in this specific section, we will use direct subscript to denote the iteration counters and use the subscript of []delimited-[][$\cdot$][ ⋅ ] to denote the block index. For example, for epoch-wise algorithm such as Algorithm 1, $[x_{s,k}]_{i}$ denotes the $i$-th block of the $(s,k)$-th iteration $x_{s,k}$. In this case, suppose the decision variable $x\in\mathbb{R}^{d}$ is separated into $m$ blocks, the kernel will take the form $h(x)=\sum_{i=1}^{m}h_{i}([x]_{i}),$ and the KC-regularity (Assumption 2.9) will naturally be extended a block separable version.
@@
A clear motivation for this modification is that when the different variable blocks $[x]_{i}$ have significant differences in magnitude, it is not realistic to require a bounded kernel condition number for the whole function. One may consider a bivariate case where $h_{1}(y)=h_{2}(y)=y^{2}/2+y^{4}/4$. Letting $[x]_{1}=0$ and $[x]_{2}\to\infty$ will cause $\kappa_{h}(\mathcal{X})\to\infty$ even for singleton set $\mathcal{X}=\{x=([x]_{1},[x]_{2})\}$. However, if we separately consider each $h_{i}$, then Proposition 2.12 implies the validity of Assumption E.1. In the special case where $h$ is element separable [3, 19], Assumption E.1 is in fact very easy to satisfy. Even if $h_{i}$ has exponentially fast growth, e.g., $h_{i}(y)=y^{2}/2+\exp\{y/R_{i}\}$, it satisfies kernel conditioning regularity with $\kappa_{h_{i}}^{\delta_{i}}=\exp\{\delta_{i}/R_{i}\}$, for $\forall\delta_{i}>0$.
@@
Under Assumption 3.1 and E.1, for any block index $i\in[m]$, denote $\mathcal{N}_{i}:=[m]\backslash\{i\}$. For any component function $f_{s}$ with $s\in[n]$, we have
@@
::=:= $L_{s}\cdot\mathrm{Diag}\Big{(}\{\nabla^{2}h_{i}([x]_{i})\}_{i=1}^{m-1}\Big{)}-\begin{bmatrix}\nabla^{2}_{1,1}f_{s}(x)&\cdots&\nabla^{2}_{1,m-1}f_{s}(x)\\ \vdots&\ddots&\vdots\\ \nabla^{2}_{m-1,1}f_{s}(x)&\cdots&\nabla^{2}_{m-1,m-1}f_{s}(x)\end{bmatrix}$
@@
where (i) is because Assumption 3.1 guarantees that $L_{s}\cdot\nabla^{2}_{\mathcal{N}_{m},\mathcal{N}_{m}}h(x)\pm\nabla^{2}_{\mathcal{N}_{m},\mathcal{N}_{m}}f_{s}(x)\succeq 0$ and $L_{s}\nabla^{2}h_{m}([x]_{m})\pm\nabla^{2}_{m,m}f(x)\succeq 0$. As a result, denoting $w=y-x$ and $\mathcal{X}^{j}:=\big{[}[x]_{j},[y]_{j}\big{]}$ for each $j\in[m]$, we can start bounding the squared difference of gradients as

### v4

@@ 2.5 A Lipschitz-like bound for gradient difference
Even with a single sample $\xi$, the above MSE will automatically go to 0 if $\|x-x_{\text{ref}}\|\to 0$. For the variance reduced methods whose design ensures this to happen, given an accurate enough estimator $\tilde{\nabla}f(x_{\text{ref}})\approx\nabla f(x_{\text{ref}})$, then $\tilde{\nabla}f(x_{\text{ref}})+\Delta_{\xi}$ will give a much more accurate estimation of $\nabla f(x)$. This forms the basic insight why variance reduction accelerates SGD under classic $L$-smoothness condition.
@@
$\mbox{(positive definiteness)}\qquad\rho(x,y)\geq 0\mbox{ and }\rho(x,y)=0\Longleftrightarrow x=y,\qquad\qquad\qquad\qquad$
$\mbox{(symmetry)}\qquad\rho(x,y)=\rho(y,x),\qquad\qquad\qquad\qquad\qquad\qquad\quad$
$\mbox{(triangle inequality)}\qquad\rho(x,y)\leq\rho(x,z)+\rho(z,y),\qquad\qquad\qquad\qquad\qquad\!\!\quad\qquad$
@@ 3.1 The general algorithm and analysis framework
1Input: Initial point $x_{1,0}$, constant $\delta$, epoch length $\tau$, step size $\eta$, interpolation factor $\gamma\in(0,1]$.
@@
3   Construct a convex set $\mathcal{X}_{s}\supseteq{B}(x_{s,0},\delta/2)$ such that $\kappa_{h}(\mathcal{X}_{s})\leq\kappa_{h}^{\delta}$ .
@@ 7      If $k\geq 1$, uniformly sample a mini-batch $\mathcal{B}_{s,k}\subseteq[n]$ with replacement, compute
$\bar{x}_{s,k+1}=\mathbf{T}_{\phi+\mathrm{id}_{\mathcal{X}_{s}}\!,h}^{\eta}(x_{s,k},v_{s,k})\qquad\mbox{and}\qquad x_{s,k+1}=(1-\gamma)x_{s,k}+\gamma\bar{x}_{s,k+1}\vskip-2.84544pt$ (18)
@@
Using the fact that $L^{2}=\frac{1}{n}\sum_{i=1}^{n}L_{i}^{2}$ and the fact that $\kappa_{h}(\mathcal{X}_{s})\leq\kappa_{h}^{\delta}$ leads to (22). ∎
@@ 3.2.2 General non-element-separable cases
For the linear consensus constraint xy=0x-y=0, the corresponding coefficient matrix is $I_{d\times d},-I_{d\times d}]$ and its condition number is 1. Then many primal-dual algorithms can achieve an $O(\kappa_{h}^{\delta}\ln\epsilon_{\mathrm{tol}}^{-1})$ complexity for finding an $\epsilon_{\mathrm{tol}}$-optimal solution, see [49]. Due to Lemma 3.9, we only need to deal with this scenario for limited times. This is also how we implement the subproblem solvers in the experiments.
@@ 3.2.3 Composition of norm cases
Computing the objective gradient of (40) gives $H^{\prime}(\alpha)-\|u\|$. Because $H(\cdot)$ is strictly convex and increasing, we know $H^{\prime}(\cdot)\geq 0$ and $H^{\prime}(\cdot)$ is increasing on $0,+\infty)$, then we know $\alpha^{*}=0$ if $H^{\prime}(0)\geq\|u\|$. Otherwise, there is a unique solution $\alpha^{*}$ s.t. $H^{\prime}(\alpha^{*})=\|u\|$. In some cases, the equation $H^{\prime}(\alpha)=\|u\|$ allows a closed form solution, then we can directly adopt it, see [5]. If no closed-form solution is available, then one can apply either Newton’s method or a binary search to find the root. Based on this observation, let us consider a few more examples.
@@
In fact, this result can be generalized to the group $\ell_{1}$$\ell_{2}$ norm that promotes group sparsity.
@@ 4.1 An adaptive step size control for BPG method
$\Psi(x_{k+1})\leq\Psi(x_{k})-\frac{\min\left\{\rho\mu\delta,\,\,3\min\big\{\frac{1}{2L},\frac{\mu\delta}{3\rho}\big\}\cdot\big\|\mathcal{D}_{\phi,h}^{\lambda_{k}}(x_{k})\big\|^{2}\right\}}{4L_{h}([x_{k},x_{k+1}])}.$
@@ 4.2 Adaptive step size control with stochastic variance reduction
$\bar{x}_{s,k+1}=\mathbf{T}_{\phi,\,h}^{\eta_{s,k}}\big(x_{s,k},v_{s,k}\big)\qquad\mbox{with}\qquad\eta_{s,k}=\min\left\{\frac{1}{2\kappa_{h}^{\delta}L},\,\frac{\mu\delta}{3\rho},\,\frac{\mu\delta}{\|v_{s,k}\|+\rho}\right\}\,$$\quad$ (46)
@@
[…]
@@
Suppose Assumptions 2.9 and 3.1 hold. For any epoch $s\geq 1$ and let $p_{s}=\frac{6q}{\pi^{2}s^{2}\tau}$ for some $q\in(0,1)$. Suppose we select $|\mathcal{B}_{s,k}|=b_{s}$ for all $0\leq k\leq\tau-1$, then
@@
In the above definition, the index $k$ runs through $1\leq k\leq\tau-1$, and the index $j$ can take value from $1\leq j\leq|\mathcal{B}_{s,k}|$ given each $k$. For each $(k,j)$ in our index range, $\xi_{k,j}^{s}$ stands for the $j$-th sample from the batch $\mathcal{B}_{s,k}\subseteq[n]$. Then by direct computation, we have $\mathcal{E}_{s,k}=\sum_{k^{\prime}=1}^{k}\sum_{j=1}^{|\mathcal{B}_{s,k^{\prime}}|}\zeta_{k^{\prime},j}^{s}$ and $\{\zeta_{k,j}^{s}\}$ forms a martingale difference sequence if the index $(k,j)$ runs in a lexicographical order. Note that
@@
Let us set the parameters according to Lemma 4.9 and 4.10. Consider the target accuracy $\epsilon\leq\max\left\{2L\kappa_{h}^{\delta}\mu\delta\rho/3,\rho^{2}\right\}$, let $\mathcal{X}_{\epsilon}:=\mathrm{Lev}_{0}\cap B(x_{1,0},R_{\epsilon})$ be a compact set, then with probability at least 1q1-q, all the following arguments hold:
@@
[…]
@@ 5 Numerical experiments
Figure 4: Experimental results for quadratic reverse problem. The raw signal vector $x_{\mathrm{true}}$ is attached at the bottom left corner of each subfigure. The “# of datapass” in x-axis refers to $\frac{\#\mbox{samples consumed}}{\mbox{full batch size }N}$.
@@ Appendix B Proof of Section 2
First, direct computation gives $\nabla^{2}h(x)=(\|x\|^{r}+\alpha)\cdot I+r\|x\|^{r-2}\cdot xx^{\top}$. For $\forall x\in\mathbb{R}^{d}$, we have $\lambda_{\max}(\nabla^{2}h(x))=(r+1)\|x\|^{r}+\alpha$ and $\lambda_{\min}(\nabla^{2}h(x))=\|x\|^{r}+\alpha$. Then for any compact set $\mathcal{X}\subseteq\mathbb{R}^{n}$ with diameter denoted by $\mathrm{diam}(\mathcal{X})=d_{\mathcal{X}}$, let $y\in\mathop{\mathrm{argmax}}_{u\in\mathcal{X}}\|u\|$ and $x\in\mathop{\mathrm{argmin}}_{u\in\mathcal{X}}\|u\|$, then
$\kappa_{h}(\mathcal{X})=\frac{\lambda_{\max}(\nabla^{2}h(y))}{\lambda_{\min}(\nabla^{2}h(x))}=\frac{(r+1)\|y\|^{r}+\alpha}{\|x\|^{r}+\alpha}\leq\frac{(r+1)(\|x\|+d_{\mathcal{X}})^{r}+\alpha}{\|x\|^{r}+\alpha}.$ (52)
@@
$\kappa_{h}(\mathcal{X})\leq(r+1)\left(1+\frac{d_{\mathcal{X}}}{\|x\|}\right)^{r}+\frac{\alpha}{\|x\|^{r}+\alpha}\leq(r+1)\left(1+\frac{1}{r}\right)^{r}+1\leq 3r+4,\quad\forall r\geq 0.$ (53)
@@
$\kappa_{h}^{\delta}\leq(r+1)\cdot\sup\left\{\frac{(t+\delta)^{r}}{t^{r}+\alpha}:t\geq 0\right\}+1.$ (54)
By direct computation, the function $\rho(t):=\frac{(t+\delta)^{r}}{t^{r}+\alpha}$ attains its maximal value at
[…]
@@
In particular, when $r>1$ and $\delta\leq\alpha^{\frac{1}{r}}/r\leq\alpha^{\frac{1}{r}}/r^{\frac{r-1}{r}}$, we have $\big(\frac{\delta^{r}}{\alpha}\big)^{\frac{1}{r-1}}\leq\frac{1}{r-1}$. Following the same logic of the last inequality in (53), we obtain $\kappa_{h}^{\delta}\leq 3r+4$ in this case. This completes the proof. ∎
@@
Let $A$ be a symmetric matrix, and let $B\succeq 0$ be positive semidefinite matrix. Then $B\preceq A\preceq B$ indicates that $\|A\|\leq\|B\|$.
@@
As $\mathcal{X}$ is convex, the line segment $x,y]\subseteq\mathcal{X}$ and hence
@@ Appendix E Extension to multi-block kernel conditioning
Finally, we will roughly discuss how one can extend KC-regularity to block separable kernels, which is considered in [3, 15, 16, 19]. In particular, we will mostly focus on the extension of Section 3, while the extension of Section 4 can be done similarly. To differentiate the meaning of the subscripts, in this specific section, we will use direct subscript to denote the iteration counters and use the subscript of $\cdot$] to denote the block index. For example, for epoch-wise algorithm such as Algorithm 1, $x_{s,k}]_{i}$ denotes the $i$-th block of the $(s,k)$-th iteration $x_{s,k}$. In this case, suppose the decision variable $x\in\mathbb{R}^{d}$ is separated into $m$ blocks, the kernel will take the form $h(x)=\sum_{i=1}^{m}h_{i}([x]_{i}),$ and the KC-regularity (Assumption 2.9) will naturally be extended a block separable version.
@@
A clear motivation for this modification is that when the different variable blocks $x]_{i}$ have significant differences in magnitude, it is not realistic to require a bounded kernel condition number for the whole function. One may consider a bivariate case where $h_{1}(y)=h_{2}(y)=y^{2}/2+y^{4}/4$. Letting $x]_{1}=0$ and $x]_{2}\to\infty$ will cause $\kappa_{h}(\mathcal{X})\to\infty$ even for singleton set $\mathcal{X}=\{x=([x]_{1},[x]_{2})\}$. However, if we separately consider each $h_{i}$, then Proposition 2.12 implies the validity of Assumption E.1. In the special case where $h$ is element separable [3, 19], Assumption E.1 is in fact very easy to satisfy. Even if $h_{i}$ has exponentially fast growth, e.g., $h_{i}(y)=y^{2}/2+\exp\{y/R_{i}\}$, it satisfies kernel conditioning regularity with $\kappa_{h_{i}}^{\delta_{i}}=\exp\{\delta_{i}/R_{i}\}$, for $\forall\delta_{i}>0$.
@@
[…]
@@
$:=$ $L_{s}\cdot\mathrm{Diag}\Big(\{\nabla^{2}h_{i}([x]_{i})\}_{i=1}^{m-1}\Big)-\begin{bmatrix}\nabla^{2}_{1,1}f_{s}(x)&\cdots&\nabla^{2}_{1,m-1}f_{s}(x)\\ \vdots&\ddots&\vdots\\ \nabla^{2}_{m-1,1}f_{s}(x)&\cdots&\nabla^{2}_{m-1,m-1}f_{s}(x)\end{bmatrix}$
@@
where (i) is because Assumption 3.1 guarantees that $L_{s}\cdot\nabla^{2}_{\mathcal{N}_{m},\mathcal{N}_{m}}h(x)\pm\nabla^{2}_{\mathcal{N}_{m},\mathcal{N}_{m}}f_{s}(x)\succeq 0$ and $L_{s}\nabla^{2}h_{m}([x]_{m})\pm\nabla^{2}_{m,m}f(x)\succeq 0$. As a result, denoting w=yxw=y-x and $\mathcal{X}^{j}:=\big[[x]_{j},[y]_{j}\big]$ for each $j\in[m]$, we can start bounding the squared difference of gradients as